# Let Occ Flow: Self-Supervised 3D Occupancy Flow Prediction

### v2

@@ B.3 Training Settings
The resolution of the input image is 512x1408 for nuScenes [42], 352x1216 for SemanticKITTI [39], and 352x1216 for KITTI-MOT [38]. For loss weight, we set $\lambda_{flow}=5\times 10^{-3},\lambda_{eik}=0.1,\lambda_{dreg}=0.1$, if present, and the weights for the edge $\lambda_{edge}$ and the LiDAR $\lambda_{lidar}$ losses are 0.02 and 0.2 respectively. During training, we adopt the AdamW optimizer with an initial learning rate 1e-4 and weight decay of 0.01. We use the multi-step learning rate scheduler with linear warming up in the first 1k iterations. We train our models with a total batch size of 8 on 8 A100 GPUs for 16 epochs. Experiments on SemanticKITTI [39] and KITTI-MOT [38] take less than one day, while experiments on nuScenes [42] finish within two days.
@@ C.1 Depth Estimation Evaluation Metrics
To demonstrate the advanced occupancy prediction quality of our approach, we employ Ray-based IoU (RayIoU[40] as our evaluation metric and re-evaluate other comparative methods using their open-source checkpoint models. Initially, we generated LiDAR rays with origins sampled from the sequence, and then compute the travel distances of each LiDAR ray before intersecting with the occupied voxels. A query ray is classified as true positive(TP) when the L1 error between the predicted depth and ground-truth depth is less than a pre-defined threshold (1m, 2m, and 4m).
@@ C.3 Occupancy Flow Prediction Evaluation Metrics
On the KITTI-MOT [38] data, due to the absence of ground-truth occupancy flow labels, we conduct the evaluation following the scene flow benchmark of KITTI dataset [38]. Specifically, we use LiDAR-projected depth and pseudo 2D optical flow cues to generate disparity and optical flow labels. And we computed end-point error of rendered and ground-truth disparity (DE) and optical flow (EPE). D1_5represent the percentage of disparity outliers with end-point error smaller than 4 pixels or 5% of the ground-truth disparity. F1_10is the percentage of optical flow outliers with end-point error smaller than 8 pixels or 10% of the ground-truth flow labels. And SF_10is the percentage of scene flow outliers (outliers in either D1_10or F1_10%). To better evaluate the scene flow in foreground regions, we also reported the foreground disparity error (DE_FG) and optical flow error (EPE_FG) by leveraging the semantic mask obtained from GroundedSAM [16].
@@
Appendix D Visualization Results
D.1 Depth Estimation
Figure 7 shows the qualitive comparison of depth estimation on the SemanticKITTI [39] validation set. The visualization results illustrate the successful prediction of detailed and accurate depth by our method compared to other rendering-based approaches.
D.2 3D Occupancy Prediction
Figure 8 shows the qualitive comparison of 3D occupancy prediction on the SemanticKITTI [39] validation set. Our method achieves precise occupancy prediction for thin structures such as poles, trees, and cyclists. Also, it provides smooth predictions for cars and road surfaces compared to other supervised [41] and self-supervised [25, 5, 6] methods.
We provided more visualization results of depth estimation, novel view depth synthesis, and 3D occupancy prediction to illustrate the superiority of our method. As is shown in Figure 9, our method exhibited strong performance across these three tasks.
D.3 Occupancy Flow Prediction
Figure 10 and Figure 11 shows the visualization results of the depth estimation and occupancy flow prediction tasks on the KITTI-MOT [38] and nuScenes [42] validation set. Our proposed method can simultaneously provide accurate 3D occupancy and occupancy flow prediction.
@@ D.3 Occupancy Flow Prediction
Appendix E Limitations and Future Work
Although we use temporal sequence input to better exploit the historical information, our model cannot completely handle the occlusion problem due to the inherent rendering-based limitation. Subsequent research could investigate long-term occupancy flow modeling and solutions to leverage the temporal sequence supervision to scale up the visible range of perspective. In addition, although our Let Occ Flow can offer accurate 3D occupancy and occupancy flow prediction, the prediction quality highly relies on the quality of optical flow cues, which serves as the supervision for the flow field. To improve the quality of 2D optical flow labels, we proposed a tracking-based flow estimation method and an unsupervised flow fine-tuning strategy. We hope the future work can pay more attention on the improvement of the flow supervision quality. Finally, our occupancy flow prediction does not explicitly enforce consistency within instances, and future work may explore to integrate instance perception into occupancy flow prediction.

### v3

@@ B.3 Training Settings
The resolution of the input image is 512x1408 for nuScenes [39], 352x1216 for SemanticKITTI [35], and 352x1216 for KITTI-MOT [37]. For loss weight, we set $\lambda_{flow}=5\times 10^{-3},\lambda_{eik}=0.1,\lambda_{dreg}=0.1$, if present, and the weights for the edge $\lambda_{edge}$ and the LiDAR $\lambda_{lidar}$ losses are 0.02 and 0.2 respectively. During training, we adopt the AdamW optimizer with an initial learning rate 1e-4 and weight decay of 0.01. We use the multi-step learning rate scheduler with linear warming up in the first 1k iterations. We train our models with a total batch size of 8 on 8 NVIDIA A100 80GB GPUs with 8 epochs for the first stage, and 16 epochs for the second stage. Experiments on SemanticKITTI [35] and KITTI-MOT [37] take less than one day, while experiments on nuScenes [39] finish within two days.
@@ B.5 Differentiable Rendering
B.6 Regularization.
Due to the explicit physical interpretation of SDF representation, regularization can be effectively applied to the volumetric SDF field to prevent over-fitting with limited supervision. Initially, an eikonal term is employed for each sample point along camera and LiDAR rays to promote the establishment of a valid SDF field:
$\mathcal{L}_{eik}=(1-\|\left\langle\nabla\mathcal{S},\bm{p}\right\rangle)\|_{2})^{2}$ (15)
where $\nabla\mathcal{S}$ denotes the pre-computed gradient field of the volumetric SDF grid. Additionally, we also utilize a Hessian loss $\mathcal{L}_{H}$ on the second-order derivatives of the SDF grid, following the approach of SelfOcc [6]. Furthermore, an edge-aware smoothness loss $\mathcal{L}_{smooth}$ is employed following monodepth2 [33] to enforce neighborhood consistency of rendered depths and flows.
@@ C.1 Depth Estimation Evaluation Metrics
To demonstrate the advanced occupancy prediction quality of our approach, we employ Ray-based IoU (RayIoU[36] as our evaluation metric and re-evaluate other comparative methods using their open-source checkpoint models. Initially, we generated LiDAR rays with origins sampled from the sequence, and then compute the travel distances of each LiDAR ray before intersecting with the occupied voxels. A query ray is classified as true positive(TP) when the L1 error between the predicted depth and ground-truth depth is less than a pre-defined threshold (1m, 2m, and 4m). We finally average over distance thresholds and compute the mean RayIoU metric.
@@ C.3 Occupancy Flow Prediction Evaluation Metrics
On the KITTI-MOT [37] data, due to the absence of ground-truth occupancy flow labels, we conduct the evaluation following the scene flow benchmark of KITTI dataset [37]. Specifically, we use LiDAR-projected depth and pseudo 2D optical flow cues to generate disparity and optical flow labels. And we computed end-point error of rendered and ground-truth disparity (DE) and optical flow (EPE). D_5represent the percentage of disparity outliers with end-point error smaller than 4 pixels or 5% of the ground-truth disparity. Fl_10is the percentage of optical flow outliers with end-point error smaller than 8 pixels or 10% of the ground-truth flow labels. And SF_10is the percentage of scene flow outliers (outliers in either D_10or Fl_10%). To better evaluate the scene flow in foreground regions, we also reported the foreground disparity error (DE_FG) and optical flow error (EPE_FG) by leveraging the semantic mask obtained from GroundedSAM [16].
@@
Appendix D Additional Ablative Studies
D.1 Flow Supervision
To demonstrate the effectiveness to introduce 2D optical flow supervision into our approach, we conducted a supplementary ablation by dropping the flow supervision.
Dropping flow supervision would result in zero flow predictions in dynamic regions. The ambiguity between camera and object motion makes the photometric loss insufficient for jointly predicting accurate geometry and motion. This challenge is even more pronounced in occupancy-level geometry and motion prediction.
As shown in Table 7, while depth-related metrics (DE, DE_FG, D_5) exhibit minimal variance, there is a notable increase in the End-Point Error of foreground objects (EPE_FG), the optical flow outlier rate(Fl_10%) and the scene flow outlier rate(SF_10%), underscoring the importance of flow supervision in our model.
Table 7: Static flow supervision ablation on KITTI-MOT [37] dataset.
[…]
2.447 3.649 2.936 9.663 0.149 0.114 0.139
$\checkmark$ 2.610 3.488 3.109 7.510 0.154 0.083 0.105
D.2 Optimization Strategy on nuScenes
To demonstrate the results of our ablations on different datasets, we conducted a supplementary ablation by employing nuScenes with $256\times 704$ image resolution and a half-epoch setting (8 epochs). This additional analysis serves to underscore the effectiveness of our method.
As demonstrated in Table 8, the ablation results on nuScenes align well with those on KITTI-MOT. The joint optimization of geometry and motion without two-stage optimization results in poor geometric performance due to the inherent ambiguities of ego and object motion. The average velocity error (mAVE) only accounts for true positive predictions, thus underestimating our full model by disregarding incorrect occupancy predictions in dynamic regions. Additionally, Our proposed dynamic disentanglement strategy enhances the performance by introducing geometric regularization of static regions and addresses the imbalance problem.
Table 8: Optimization strategy ablation on nuScenes [39] dataset. Since the absence of two-stage optimization brings instability during training, we use extra LiDAR ray supervision here.
Two Stage Optimization Dynamic Disentanglement 3D Occupancy & Occupancy Flow
[…]
10.57 21.05 41.15 24.26 1.36
$\checkmark$ 16.44 30.53 50.51 32.49 1.59
$\checkmark$ $\checkmark$ 22.95 36.82 54.00 37.92 1.56
Appendix E Visualization Results
E.1 Depth Estimation
Figure 7 shows the qualitive comparison of depth estimation on the SemanticKITTI [35] validation set. The visualization results illustrate the successful prediction of detailed and accurate depth by our method compared to other rendering-based approaches due to our effective temporal fusion module and optical flow supervision.
E.2 3D Occupancy Prediction
Figure 8 shows the qualitive comparison of 3D occupancy prediction on the SemanticKITTI [35] validation set. 3D occupancy are color-coded according to the height of the voxels. Our method achieves precise occupancy prediction for thin structures such as poles, trees, and cyclists. Also, it provides smooth predictions for cars and road surfaces compared to other supervised [38] and self-supervised [5, 6, 25] methods.
We provided more visualization results of depth estimation, novel view depth synthesis, and 3D occupancy prediction to illustrate the superiority of our method. As is shown in Figure 9, our method exhibited strong performance across these three tasks, which is due to the accurate prediction of the visible surfaces. In addition, our method with effective temporal fusion can provide occlusion perception beyond the visible surface, thus generating high-quality novel-view depth estimations and global-view occupancy predictions.
E.3 Occupancy Flow Prediction
Figure 10 and Figure 11 shows the visualization results of the depth estimation and occupancy flow prediction tasks on the KITTI-MOT [37] and nuScenes [39] validation set. For flow visualization, the color and brightness represent the motion direction and magnitude respectively. Our proposed method can simultaneously provide accurate 3D occupancy and occupancy flow prediction.